A whip cream container that is 0.0C is placed into a campfire and its temperature is raised to 300 C, what was the initial pressure in the container if the final pressure is 19 atmospheres?
Answer #1

We know that equation,

P1/T1 = P2/T2 thus,

P1 = T1P2/T2 where,

P1 = initial pressure =?

P2 = final pressure = 19 atm

T1 = initial tempreture = 00C = 273.15 K

T2 = final tempreture = 3000C = 573.15K

substitute the value in above equation

P1 = 273.15 \times19 / 573.15 = 9.055 atm

initial pressure in the container = 9.055 atm

Note - I have converted tempreture in kelvin because 00C mulply by 19 give answer 0 so only for mathematical calculation i have converted in kelvin otherwise not need to convert in kelvin.